I need help with this problem:
log(x-7) - log(8x-50) = log(1/x)
I need the exact solution and a two decimal place approximation when appropriate.
The book says that the answer is 10 but I'm not arriving at that answer when doing the problem.
Thank you!

Answer by ewatrrr(24785)

x(x-7)/(8x-50) = 1
x^2 -15x + 50 = 0
(x-10)(x-5) = 0, x = 10 |Note: x = 5 is an extraneous root (x-7)CANNOT be < 0